# OpenAD: Open-World Autonomous Driving Benchmark for 3D Object Detection

### v1

@@ 5.1 Vision-Centric 3D Open-ended Object Detection
Due to the limited scale of existing 3D perception data, it is challenging to directly train a vision-based 3D open-world perception model. We utilize existing 2D models with strong generalization capabilities to address this issue and propose a vision-centric baseline for 3D open-world perception.
As illustrated in Figure 4, an arbitrary existing 2D open-world object detection method is initially employed to obtain 2D bounding boxes and their corresponding semantic labels. Simultaneously, the image feature maps generated by the image encoder of the 2D model are cached. Subsequently, a 2D-to-3D Bbox Converter, which combines multiple features and a few trainable parameters, is introduced to transform 2D boxes into 3D boxes.
Figure 4: The 3D open-world object detection baseline we proposed. Based on the existing 2D open-world models and depth estimation models, we train a 2D-to-3D Bbox Converter. This module extracts object features through pseudo-point clouds and convolutional dual-branch architecture to predict the 3D bounding boxes of objects.
Specifically, we use existing depth estimation models, such as ZoeDepth [4], DepthAnything [65], and UniDepth [49], to obtain the depth map of the cropped image by the 2D box. We also include an optional branch that utilizes LiDAR point clouds and a linear fitting function to refine the depth map by projecting point clouds onto the image. Simultaneously, to eliminate regions within the 2D bounding box that do not belong to the foreground object, we utilize Segment Anything Model [33] (SAM) to segment the object with the 2D box as the prompt, yielding a segmentation mask. After that, we can construct pseudo point clouds for the segmentation mask with its pixel coordinates, depth map, and camera parameters. We project the pseudo point cloud onto the feature map and depth map, and features are assigned to each point through interpolation. Then, we adopt PointNet [51] to extract the feature $f_{p}$ of the pseudo point clouds. Meanwhile, the depth map and feature map within the 2D bounding box are concatenated along the channel dimension, and its feature $f_{c}$ is derived through convolution and global pooling. Finally, we utilize an MLP to predict the object’s 3D bounding box with the concatenated features of $f_{p}$ and $f_{c}$.
In this baseline, only a few parameters in the 2D-to-3D Bbox Converter are trainable. Thus, the training cost is low. In addition, during the training, each 3D object serves as a data point for this baseline, allowing for the straightforward construction of multi-domain dataset training.
@@ 5.2 General and Specialized Models Fusion
In experiments, we have found that existing open-world methods or general models are inferior to close-set methods or specialized models in handling objects belonging to common categories, but they exhibit stronger domain generalization capabilities and the ability to deal with corner cases. That is to say, existing general and specialized models complement each other. Hence, we leverage their strengths and propose a fusion baseline by combining the prediction results from the two types of models. Specifically, we align the confidence scores of the two types of models and perform non-maximum suppression (NMS) with dual thresholds, i.e., IoU and semantic similarity, to filter duplicates.
@@ 6 Experiments
Table 3: Evaluation of 3D open-world methods (top), specialized methods (middle), and ensemble methods (bottom) on OpenAD benchmark. ARnusc refers to scenes derived from nuScenes in OpenAD, with ARseen denoting object categories observed in the nuScenes training set. All methods are trained on nuScenes.
@@
OpenAD-G LC GenerateU 9.02 23.32 0.970 0.521 19.79 7.14 25.78 10.15
OpenAD-Y LC YOLOWorld 9.43 25.17 0.872 0.535 25.54 13.83 31.31 9.84
@@
OpenAD-Ens LC OpenAD-Y + BEVFusion 16.34 44.16 0.792 0.469 62.14 13.83 35.41 9.84
6.1 Evaluation Details
For 2D open-vocabulary methods, which need a predefined object category list from users as additional inputs to detect corresponding objects, we take the union of the categories from five datasets and incorporate two additional open-vocabulary queries, i.e., “object that affects traffic” and “others”, into it. We adopt OWLv2-CLIP-L/14-ST+FT, YOLO-Worldv2-XL, GLIP-L, and GroundingDINO-B for OWL-ViT v2 [47], YOLO-World v2 [15], GLIP [35], and GroundingDino [42], respectively.
2D open-ended methods can directly provide bounding boxes and corresponding natural language descriptions, enabling direct evaluation for OpenAD. We employ the “vg-grit5m” version for GenerateU [16].
For the 3D open-ended baselines we proposed, the 2D-to-3D Bbox Converter is trained on nuScenes. We use GenerateU [16] and YOLO-World [15] as the 2D detector, Depth Anything [65] as the depth estimation model, and SAM [33] as the segmentation model. All these 2D models are frozen without any fine-tuning.
@@ 6.2 Main Results
As shown in Tables 2 and 3, we conduct evaluations on various 2D and 3D object detection models, including 2D and 3D open-world models, specialized models, and our baselines.
The results show that current open-world models, irrespective of being 2D or 3D detectors, tend to predict objects unrelated to driving (such as the sky) or to make repeated predictions for different parts of the same object, resulting in low precision and AP. Nevertheless, these models demonstrate good domain generalization and open-vocabulary capabilities, which are lacking in current specialized models. Note that our proposed ensemble baselines can effectively combine the advantages of open-world and specialized models, achieving favorable performance in both seen and unseen domains and categories. In addition, in Table 3, our proposed vision-centric baseline for 3D open-world object detection leverages the capabilities of 2D open-world models. Specifically, by harnessing the open-world capabilities of Yolo-world v2, our method obtains 0.58 AP and 6.2 AR improvement compared to Find n’ Propagate.
Moreover, we observed that the issue of overfitting is more pronounced for 3D object detection models on datasets such as nuScenes. Some models perform superior in-domain benchmarks but show worse domain generalization ability. For instance, SparseBEV, compared to methods based on Lift-Splat-Shot, achieves impressive in-domain results, with its in-domain AR even surpassing those of LiDAR-based methods. However, SparseBEV’s domain generalization capability is relatively poor. Models with increased parameters by enlarging the backbone, including BEVStereo and SparseBEV, show more severe overfitting issues. These results reveal the limitations of in-domain benchmarks like nuScenes. In contrast, augmenting the parameter count through utilizing BEVFormer v2 or HENet simultaneously enhances both in-domain and out-domain Recall, indicating an inherent improvement in the methodology. Therefore, even for specialized models trained on a single domain, evaluating them on OpenAD benchmarks remains meaningful.
@@ 6.3 Ablations of Proposed Baselines
We conduct ablation studies for the proposed baselines, as shown in Table 4. We find that additional Pseudo Point Cloud inputs bring 9.9 mAR. In addition, replacing MLP with unlearnable PCA methods decreases the performance by a large margin, from 45.1 mAR to 27.3 mAR. These results show that the simple MLP can learn to complete the boundaries of objects from the datasets and predict more accurate 3D boxes.
Table 4: Ablation of 2D-to-3D Bbox Converter. This module is trained using the 2D-3D annotation pairs from the nuScenes training set and tested on the 2D-3D annotation pairs from OpenAD.
Conv Pseudo Point Cloud Bbox Decoding mAR
MLP 36.8
PCA for Oriented Bounding Box 27.3
MLP 45.1
MLP 46.7
@@ 7 Conclusion
In this paper, we introduce OpenAD, the first open-world autonomous driving benchmark for 3D object detection. OpenAD is built on a corner case discovery and annotation pipeline that integrates with a multimodal large language model. The pipeline aligns five autonomous driving perception datasets in format and annotates corner case objects for 2000 scenarios. In addition, we devise evaluation methodologies and analyze the strengths and weaknesses of existing open-world perception models and autonomous driving specialized models. Moreover, addressing the challenge of training 3D open-world models, we proposed a baseline method for 3D open-world perception by combining 2D open-world models. Furthermore, we introduce a fusion baseline approach to leverage the advantages of open-world models and specialized models.
@@
We hope that OpenAD can help develop open-world perception models that surpass specialized models, whether in the same domain or across domains, and whether for semantic categories that have been seen or unseen.

### v2

@@ 5.1 Vision-Centric 3D Open-ended Object Detection
Figure 4: The 3D open-world object detection framework we proposed. After obtaining 2D proposals from any frozen open-world 2D object detection model, we train a 2D-to-3D BBox Converter to predict 3D bounding boxes. The converter has a dual-branch architecture, which extracts pseudo-point features and convolutional features. It is lightweight and easy to train.
Due to the limited scale of existing 3D perception data, it is challenging to train a vision-based 3D open-world perception model directly. To address this issue, we propose a vision-centric framework for 3D open-world perception, as illustrated in Figure 4. An existing 2D open-world object detection model is first employed to generate 2D proposals and their corresponding semantic labels. Subsequently, a 2D-to-3D BBox Converter is introduced, which combines multiple features and a few trainable parameters, to transform 2D proposals into 3D boxes.
Specifically, for each 2D proposal, we employ a Partial Encoder composed of multi-layer convolutional networks to extract partial features. This is followed by a Depth Net constructed with MLPs to predict depth at each grid, ultimately generating a depth map. We also include an optional branch that utilizes LiDAR point clouds and a linear fitting function to refine the depth map by projecting point clouds onto the image. A 3D point coordinate can be obtained from the coordinates of each grid, camera parameters, and depth. Pseudo point clouds with features are obtained in this way. We project the pseudo point clouds onto the feature map and depth map, and features are assigned to each point through interpolation. Then, we adopt PointNet [49] to extract the feature $f_{p}$ of the pseudo point clouds. Meanwhile, the depth and feature maps within the 2D bounding box are concatenated along the channel dimension, and its feature $f_{c}$ is derived through convolution and global pooling. Finally, we utilize an MLP to predict the object’s 3D bounding box with the concatenated features of $f_{p}$ and $f_{c}$.
Only a few parameters in the 2D-to-3D Bbox Converter are trainable in this baseline. Thus, the training cost is low. In addition, during the training, each 3D object serves as a data point for this baseline, allowing for the straightforward construction of multi-domain dataset training.
@@ 5.2 General and Specialized Models Fusion
In experiments, we find that existing open-world methods or general models are inferior to close-set methods or specialized models in handling objects belonging to common categories, but they exhibit stronger domain generalization capabilities and the ability to deal with corner cases. That is to say, existing general and specialized models complement each other. Hence, we utilize their strengths and propose a fusion baseline by combining the prediction results from the two types of models. Specifically, we align the confidence scores of the two types of models and perform non-maximum suppression (NMS) with dual thresholds, i.e., IoU and semantic similarity, to filter duplicates.
@@ 6 Experiments
6.1 Implementation Details
Open-ended 2D methods can directly provide bounding boxes and the corresponding natural language descriptions, enabling direct evaluation of OpenAD. For 2D open-vocabulary methods, which need a predefined object category list from users as additional inputs to detect corresponding objects, we take the union of the common categories from five datasets and incorporate two additional open-vocabulary queries, i.e., “object that affects traffic” and “others”, into it. For all open-world 2D models, we selected their best-performing open-source model and conducted zero-shot evaluation on OpenAD.
For the baseline method we proposed, the 2D-to-3D Converter is trained on nuScenes (10 common classes only) for 10 epochs, over 20 hours using 4 A40 GPUs. We use GenerateU [15] and YOLO-World [14] to generate the 2D proposals, respectively. They are frozen without any fine-tuning.
@@ 6.1 Implementation Details
Table 3: Evaluation of 3D open-world methods (top), specialized methods (middle), and ensemble methods (bottom) on OpenAD benchmark. ARnusc refers to scenes derived from nuScenes in OpenAD, with ARseen denoting object categories observed in the nuScenes training set. All methods are trained on nuScenes training set.
@@
OpenAD-G LC GenerateU 15.14 34.46 1.056 0.649 14.54 11.15 26.48 16.95
OpenAD-Y LC YOLOWorld 15.54 36.07 1.063 0.646 29.99 12.73 25.88 14.17
@@
OpenAD-Ens LC OpenAD-Y + BEVFusion 16.22 47.12 0.851 0.511 62.69 12.05 35.62 13.60
OpenAD-Ens LC OpenAD-G + BEVFusion 16.30 48.25 0.858 0.520 64.84 10.59 39.11 16.85
@@ 6.2 Main Results
Tables 2 and 3 show the evaluation results on 2D and 3D object detection models, including 2D and 3D open-world models, specialized models, and our baselines. The results show that current open-world models, irrespective of whether they are 2D or 3D detectors, tend to predict objects unrelated to driving (such as the sky) or to make repeated predictions for different parts of the same object, resulting in low precision and AP. However, these models demonstrate good domain generalization and open-vocabulary capabilities, which are lacking in current specialized models. Note that our proposed ensemble baselines can effectively combine the advantages of open-world and specialized models, achieving favorable performance in both seen and unseen domains and categories. In addition, in Table 3, our proposed vision-centric baseline for 3D open-world object detection utilizes the capabilities of 2D open-world models. Specifically, by harnessing the open-world capabilities of Yolo-world v2, our method obtains 6.7 AP and 17.1 AR improvement compared to Find n’ Propagate.
Moreover, we observed that the issue of overfitting is more pronounced for 3D object detection models on datasets such as nuScenes. Some models perform well in-domain benchmarks but show worse domain generalization ability. For instance, SparseBEV, compared to methods based on Lift-Splat-Shot, achieves impressive in-domain results, with its in-domain AR even surpassing those of LiDAR-based methods. However, SparseBEV’s domain generalization capability is relatively poor. Models with increased parameters by enlarging the backbone, including BEVStereo and SparseBEV, show more severe overfitting issues. These results reveal the limitations of in-domain benchmarks such as nuScenes. In contrast, increasing parameters through utilizing BEVFormer v2 or HENet simultaneously enhances both in-domain and out-domain Recall, indicating an inherent improvement in the methodology. Therefore, even for specialized models trained on a single domain, evaluating them on OpenAD benchmarks remains meaningful.
@@
The proposed converter enables convenient cross-dataset training. As detailed in Appendix B, AP and AR of our proposed method can be further improved by training on multiple datasets and combining 2D segmentation models.
@@ 6.3 Ablations of Proposed Baselines
We conduct ablation studies for the proposed baselines, as shown in Table 4. When decoding bounding boxes solely from either the pseudo point features $f_{p}$ or the convolutional features $f_{c}$, performance drops, demonstrating the effectiveness of our proposed dual-branch architecture. In addition, replacing MLP with unlearnable PCA methods decreases the performance by a large margin, from 23.32 AR to 13.63 mAR. These results show that the simple MLP can learn to complete the boundaries of objects from the datasets and predict more accurate 3D boxes. In initial experiments, we employed a frozen Depth Anything [63] model to obtain depth estimations. Subsequent experimental results reveal that using a lightweight trainable depth network can enhance the converter’s performance.
Table 4: Ablation of 2D-to-3D BBox Converter. This module is trained on nuScenes training set and tested on OpenAD. The 2D proposals are generated by GenerateU [15].
Conv & AvgPool Pseudo Point Clouds Depth Bbox Decoding AP AR
Frozen Depth Anything MLP 7.90 18.38
Frozen Depth Anything PCA for Oriented Bounding Box 7.61 13.63
Frozen Depth Anything MLP 8.97 22.02
Frozen Depth Anything MLP 9.02 23.32
Trainable Depth Net MLP 15.14 34.46
7 Limitations
OpenAD exclusively supports 2D and 3D object detection tasks, with all re-annotated data allocated for benchmarking purposes. In the future, we will incorporate evaluations for additional open-world perception tasks, such as occupancy prediction, while expanding data to enhance scope and diversity.
@@ 8 Conclusion
In this paper, we introduce OpenAD, the first open-world autonomous driving benchmark for 3D object detection. OpenAD is built upon a corner case discovery and annotation pipeline that is integrated with a multimodal large language model. The pipeline aligns five autonomous driving perception datasets in format and annotates corner case objects for 2000 scenarios. In addition, we devise evaluation methodologies and analyze the strengths and weaknesses of existing open-world perception models and specialized autonomous driving perception models. Moreover, to address the challenge of training 3D open-world models, we propose a novel framework for 3D open-world perception, which is lightweight and easy to train. Furthermore, we introduce a fusion baseline approach to take advantage of open-world and specialized models.